Circular No. 7038

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1998 U3 E. Meyer (Linz, Austria) and J. Jahn (Bodenteich, Germany) report

the discovery of a comet by Michael Jäger on 16- and 9-min Technical Pan film exposures. Observers generally report a 1′–2′ coma with condensation and a tail several arcmin long in p.a. 275^{o}–281^{o}. The precise CCD positions below were reported following requests from the Bureau:

1998 UT  α_{2000}  δ_{2000}  m_{1} Observer
Oct. 23.93  6^{h}43.^{m}3  +41^{o}31′  12.5 Jäger
24.85103 6 44 08.^{s}66 +41 32 51′′.0 14.1 Meyer
24.85591 6 44 08.98 +41 32 51.1 12  Pravec
24.89547 6 44 11.39 +41 32 53.1 13.0 Tichá
24.95534 6 44 14.95 +41 32 55.9 16  Mikuž
24.98812 6 44 16.86 +41 32 57.8 13.3 Kornos
25.18361 6 44 28.40 +41 33 04.6  Griffin
25.35451 6 44 38.14 +41 33 12.4  Kowalski
25.35865 6 44 38.54 +41 33 11.7 13.9 Bell
25.54384 6 44 49.63 +41 33 16.8 13.3 Asami
25.55625 6 44 50.39 +41 33 17.0 12.9 Abe
25.58280 6 44 51.92 +41 33 17.9 12.8 Kojima
25.58866 6 44 52.26 +41 33 18.2  Aoki
25.68134 6 44 57.69 +41 33 23.5 13.1 Sugie
25.70081 6 44 58.73 +41 33 23.2 12.9 Ikari

M. Jäger (east of Vienna). 0.25-m f/2.8 Schmidt camera. E. Meyer and H. Raab (Linz). 0.3-m reflector. P. Pravec (Ondřejov). 0.65-m reflector. J. Tichá and M. Tichý (Kleť). 0.57-m reflector. H. Mikuž and S. Maticic (Crni Vrh). 0.36-m reflector + V filter. L. Kornos and P. Kolény (Modra). 0.6-m reflector. I. P. Griffin (Cocoa). 0.3-m reflector. R. Kowalski (Zephyrhills). 0.18-m reflector. G. Bell and G. Hug (Eskridge). 0.30-m reflector. A. Asami (Hadano). Communicated by S. Nakano, Sumoto, Japan. H. Abe (Yatsuka). Communicated by Nakano. T. Kojima (Chiyoda). Communicated by Nakano. M. Aoki (Toyama). Communicated by Nakano. A. Sugie (Dynic Astronomical Observatory). Communicated by Nakano. Y. Ikari (Moriyama). Communiated by Nakano.

1998 October 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green